Two airplanes leave the airport at the same time, flying in opposite directions. One plane flies 100 miles per hour faster than the other. They fly for 4 hours and end up 2,000 miles apart. How fast do they each fly?

Answer by mananth(16946)

Plane II x mph
Plane I x+100
..
They are going away from each other.

So add up their speed.

combined speed = x+x+100
(2x+100 )
Time = 4 hours
Distance =2000 miles
Distance = speed * time
(2x+100 )*4= 2000
8x+400 =2000
8x=2000 -400
8x=1600
/8
x=200 mph Plane II
Plane I 200+100 =300 mph